OpenStudy (anonymous):

True or False: -3^2 = -9

OpenStudy (anonymous):

false

OpenStudy (aroub):

true

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

This is true since -3^2 = -3*3 = -9 zbay is saying false because he's probably thinking (-3)^2

OpenStudy (anonymous):

if its (-3)^2 then its false if its -(3)^2 then its true

OpenStudy (hima):

good answer mathematicalpro